# Arbitrary Precision Printed Ternary Neural Networks with Holistic Evolutionary Approximation

### v1

@@ I Introduction
In this work, we focus on designing digital approximate printed neural networks using Electrolyte-Gated FET (EGFET) technology [9], which offers good mobility characteristics and enables operation at low supply voltages [25], making it ideal for printed battery-powered applications. Recognizing the prohibitive cost of printed multipliers [16, 17], the state of the art has shifted to multiplier-less networks, such as power-of-2 quantized Multilayer Perceptrons (MLPs) [18, 19] or Ternary Neural Networks (TNNs) [20]. This allows bespoke circuit implementations to use only adders/subtractors and re-wiring for shift operations. However, [16, 17, 18, 19] require higher precision inputs to maintain acceptable accuracy, overlooking the often dominant overheads of necessary analog-to-digital converters (ADCs) to process sensor data [26, 27]. Using 11-bit inputs and simpler analog-to-binary converters (ABCs), in [20], we avoided interfacing costs. Existing approximate printed neural networks offer non-favorable accuracy-area trade-offs. The MLPs in [16, 17] offer low accuracy loss (e.g., 11-22%) but yield mainly modest hardware savings, with poor scalability for larger losses, while the multiplier-less 11-bit TNNs [20] and power-of-2 MLPs [18, 19] often suffer from high accuracy loss (e.g., >5>5%).
@@ III-A Printed ADCs
Flash ADCs (Fig. 2b) are ideal for printed applications due to their regularity and simplicity. They operate by comparing the input voltage to multiple reference voltages using an array of analog comparators. The comparator, i.e., the basic building block of the Flash ADC, is presented in Fig. 2a. This parallel, regular structure allows easy implementation. The comparators’ output represents the digital value in thermometer code, which is then converted to a binary encoding by counting the number of ‘1’s with a digital parallel counter. An NN-bit ADC requires $2^{N}-1$ analog comparators.
SAR ADCs (Fig. 2c) sequentially compare the input signal to reference levels, adjusting bit-by-bit, from most to least significant bits. They require only one analog comparator, but also need a Digital-to-Analog Converter (DAC) and more complex digital logic (SAR logic) to store intermediate results and set reference values. Moreover, SAR ADCs require registers that are highly area-inefficient in EGFET. Hence, in SAR logic we use only an NN-bit register for the output and a $\log_{2}N$ counter, with all other signals generated through counter multiplexing.
@@ III-B Arbitrary Precision Bespoke TNN Circuit
Focusing on area efficiency, a primary constraint in printed circuits [14], we use a single hidden layer with input precision ranging from 11 to 44 bits. Low-bit inputs allow for significantly smaller and less power-intensive ADCs, while also reducing the input precision requirements of the classifier (i.e., TNN). Throughout recent research efforts in ML circuits for printed electronics [16, 17, 27, 33], low input precision (i.e., below 4 bits) has been commonly used since target domains have low performance and precision needs, enabling printed circuits with acceptable area and power overheads [23].
@@
where $w_{i}$ are the neuron weights, $x_{i}$ are the input activations, and $n$ is the number of inputs. This reduces the precision required to represent $y$, with 1-1 being encoded as 0 while 11 remains 11. Note that (1) applies to any input precision, including 11-bit inputs.
@@
where $w_{j}$ are the neurons’ weights and $y_{j}$ are the outputs of the hidden layer. However, considering the aforementioned encoding for $y_{j}$, negation is achieved with a single NOT gate ($0\rightarrow 1$ and $1\rightarrow 0$). In addition, due to the bespoke implementation, if a weight is 1-1, the corresponding activation is inverted before added to the respective popcount, whereas if the weight is 0, the corresponding connection is removed. The product of a zero weight by the respective activation is always 0 (numerical value). The (numerical) 0 is in the middle of 1-1 and 11 and, thus, in our encoding 0 should be mapped to $\frac{1}{2}$ (i.e., $\frac{-1+1}{2}\rightarrow\frac{0+1}{2}$). Therefore, a $\frac{1}{2}$ is missing from the result for each removed connection due to zero weight. Hence, the output neuron can be implemented by a simple popcount unit (the sum in (2) accumulates only zeros or ones) and to obtain the correct output, a constant correction term $\frac{Z}{2}$ must be added to the popcount result, where $Z$ equals the number of zero weights in that neuron. Instead, we right-shift the popcount result $P$ and add $Z$: $2P+Z$. Note that $Z$ is a constant known post-training. Since the activation function of the output neuron is argmax, $2P+Z$ gives the same result as $P+\frac{Z}{2}$ without any rounding errors from integer division. As a result, the output neuron computes the following:
@@ IV-A2 LTG Error Analysis
If $\mathrm{LTG}_{\boldsymbol{w}}$ and $\mathrm{\hat{LTG}}_{\boldsymbol{w}}$ differ, it means that the two functions produced outputs of opposite signs–that is, one computed a positive weighted sum and the other a negative one. However, since $\mathrm{\hat{LTG}}_{\boldsymbol{w}}$ is implemented as a black-box function via CGP, we only observe its binary output (0 or 11), without any direct insight into how closely it approximated the actual weighted sum. To capture the severity of such mismatches, we define an error metric $D(\boldsymbol{x})$ that quantifies the difference/distance between the weighted sum of $\mathrm{LTG}_{\boldsymbol{w}}$ and the decision threshold at 0, using the exact LTG as the reference value. This provides a meaningful measure of how far the approximate function was from producing the correct output, assuming that $\mathrm{\hat{LTG}}_{\boldsymbol{w}}$ flipped the sign of the sum. An offset of 11 is added to handle cases where the sum is zero, but $\mathrm{\hat{LTG}}_{\boldsymbol{w}}$ wrongly produced a negative sign.
@@ V Results and Analysis
TNN Training: Exact TNN models are used as the baseline for accuracy comparisons. Input features are normalized within $[0,1]$, datasets follow a $70\%/30\%$ split for training/testing and TNNs are trained using QKeras [41] and its quantized_bits functionality. We employ the Adam optimizer for up to $30$ epochs of training (with early stopping). The learning rate for each network is obtained through Bayesian optimization, aiming to maximize the model’s inference accuracy with a maximum of $100$ attempts. Learning rate parameters are selected from the range of $0.001$ to $0.01$. Hidden neurons are selected via grid search within the $[1,50]$ range. For each dataset, we train a TNN for input precisions from 11 to 44 bits and, for each precision, select the TNN with the smallest hidden-layer size that achieves accuracy closest to the exact MLP [14].
@@ V-A1 Popcount Circuits
We generate approximate popcount circuits with the following constraints and configurations. The error limits $\tau_{mae}$ and $\tau_{wcae}$ are logarithmically distributed from $0.1$ to $0.5\cdot 2^{g}$ and from 11 to $0.5\cdot 2^{m}$, with $g=\lceil\log_{2}m\rceil$, resulting in 2,0902,\!090 approximate circuits. We also set CGP search termination criteria to $30$, $60$, and $300$ minutes for popcount sizes $m<16$, $m<32$, and $m<60$, respectively. Using BDD evaluation, we achieve average speeds of 7,7597,\!759 and 1,3621,\!362 evaluations per second for $16$-bit and $32$-bit popcounts. For $60$-bit popcount circuits, this number falls to $25$ evaluations per second. To account for this, we extend the time limit, thereby increasing the chances of identifying solutions with better area-accuracy trade-offs.
@@
A visualization of the obtained results for three popcount circuit sizes is provided in Fig. 6. The featured circuits are synthesized and evaluated based on the error metrics and methodology described above. For easier comparison, area results are normalized relative to the exact popcount circuit of the same size. Our results indicate that as the error limit increases, CGP optimization identifies increasingly area-efficient solutions. Additionally, we assess the efficacy of our approximation approach through a comparison of our popcount circuits with variants derived using the truncation approximation approach, which has previously been applied in a wide range of approximate circuits, including MLPs [42]. As shown in Fig. 6, our approximation achieves significantly superior trade-offs compared to truncation, and delivers about 22x area reduction for $\varepsilon_{mae}$ of only $0.5$, $1.1$, and $1.9$ for 88-bit, $16$-bit, and $47$-bit popcounts, respectively. Similar results to those shown in Fig. 6 are observed for all popcount circuit sizes.
@@ V-B Approximate TNN Evaluation
Fig. 9 shows the results of our multi-objective optimization for the 22-bit Pendigits dataset using various approximate LTG libraries, each evolved with a distinct error metric. Libraries containing only Pareto-optimal LTG units (“pareto”) yield superior outcomes. For high accuracy, LTG designs optimized for the $\varepsilon_{wcde}$ metric are better, while those evolved for the $\varepsilon_{mde}$ metric are preferable for maximizing area gains. Similar findings to Fig. 9 apply to other datasets. Hence, for each TNN, the required LTG circuits are approximated against both $\varepsilon_{mde}$ and $\varepsilon_{wced}$ to populate our library. In Fig. 10, we further investigate the role of our $D$ error metric in approximating LTG units for achieving area-efficient and high-accuracy approximate TNNs. Fig. 10 presents a correlation analysis between $\varepsilon_{mde}$ and classification accuracy, when all TNN hidden neurons are replaced with approximate LTGs of identical $\varepsilon_{mde}$, for a wide range of error values. Two TNNs are used in this analysis: our most complex one (Pendigits) and a relatively small one (RedWine). A strong non-linear correlation between our $\varepsilon_{mde}$ metric and the classification accuracy can be observed, especially as the network’s knowledge capacity increases–i.e., when moving from smaller (RedWine) to larger (Pendigits) networks, or from lower (11-bit) to higher (44-bit) input precision. Fig. 10 validates that we obtain the expected correlation between neural network inference accuracy and approximate arithmetic units, and confirms that our $\varepsilon_{mde}$ metric is suitable for effectively guiding approximation exploration toward high-accuracy regions. Specifically, as $\varepsilon_{mde}$ (and thus area savings) increases, there exists a broad region where minimal accuracy degradation is observed at the TNN level, maximizing, thus, the area efficiency of our approximate TNNs.
@@
Fig. 11 shows the accuracy-area trade-off for all explored solutions during our multi-objective optimization, with the Pareto fronts (per input precision) annotated by lines. As Fig. 11 includes 15,00015,\!000 designs, synthesizing all is impractical, so we report the estimated area using our precise surrogate model. Post-synthesis results for key points are presented in Table II. Note that the area estimation in Fig. 11 reflects only the TNN classifier area, excluding any interfacing costs. In Fig. 11, we include only the input precisions that result in at least one Pareto-optimal point for each dataset, since precisions that lead to only dominated solutions (i.e., with higher area and lower accuracy) are suboptimal in our evolutionary-based optimization. As shown, our framework consistently delivers a smooth accuracy-area trade-off. Supporting arbitrary input precision allows us to effectively populate the accuracy-area Pareto front, as it is primarily composed of designs with varying precisions, with only one precision, especially the low cost 11 bit, rarely dominating.
@@
Two cases require further discussion. For Pendigits with 11-bit input precision, our framework maintains a smooth accuracy-area trade-off despite a significant initial accuracy loss w.r.t the exact 11-bit TNN. Pendigits, being the most complex dataset examined, becomes highly sensitive to approximation with only 11-bit inputs. This is evident from the 11-bit exact Pendigits TNN achieving just 7676% accuracy, while the target MLP reaches 9494[14]. This further highlights that it is mandatory to support arbitrary input precision, as 11-bit precision only (e.g., [20]) may be insufficient in many cases.
@@
For Arrhythmia, we present results only for 11-bit precision, even though the 11-bit exact TNN exhibits a 22% accuracy loss compared to the respective exact MLP, while the 22-bit exact TNN features no accuracy loss. Our framework uses CGP to evolve exact LTG circuits into approximate variants and supports exact design with up to $90$ input bits. The 22-bit Arrhythmia LTG circuits feature over $100$ inputs of 22 bits, totaling more than $200$ input bits, and our CGP algorithm could not find suitable approximations within a reasonable timeframe. However, considering datasets like Arrhythmia, with more than $250$ input features, might be unrealistic for printed applications, particularly when factoring in ADC costs. Nevertheless, our framework still provides a strong trade-off, delivering 8686x lower area than the printed exact MLP [14], for only 55% accuracy loss (see Table II). In addition, our framework can be combined with pruning—which is not considered in this work to ensure fairness in comparisons—to reduce inputs to LTGs and potentially mitigate such issues. Alternatively, we can follow conventional hierarchical approaches for the generation of approximate LTGs with more than $200$ input bits. This involves breaking the LTG down to components (two or more adders and one comparator) and following a typical procedure for approximating dataflows [44, 45], e.g., creating a library of approximate components and then combining them through multi-objective optimization (e.g., NSGA-II). Such approaches might lead to some optimality loss but can address scalability issues in LTG approximation.

### v3

@@ I Introduction
In this work, we focus on designing digital approximate printed neural networks using Electrolyte-Gated FET (EGFET) technology [9], which offers good mobility characteristics and enables operation at low supply voltages [25], making it ideal for printed battery-powered applications. Recognizing the prohibitive cost of printed multipliers [16, 17], the state of the art has shifted to multiplier-less networks, such as power-of-2 quantized Multilayer Perceptrons (MLPs) [18, 19] or Ternary Neural Networks (TNNs) [20]. This allows bespoke circuit implementations to use only adders/subtractors and re-wiring for shift operations. However, [16, 17, 18, 19] require higher precision inputs to maintain acceptable accuracy, overlooking the often dominant overheads of necessary analog-to-digital converters (ADCs) to process sensor data [26, 27]. Using $1$-bit inputs and simpler analog-to-binary converters (ABCs), in [20], we avoided interfacing costs. Existing approximate printed neural networks offer non-favorable accuracy-area trade-offs. The MLPs in [16, 17] offer low accuracy loss (e.g., 11-22%) but yield mainly modest hardware savings, with poor scalability for larger losses, while the multiplier-less $1$-bit TNNs [20] and power-of-2 MLPs [18, 19] often suffer from high accuracy loss (e.g., >5>5%).
@@ III-A Printed ADCs
Flash ADCs (Fig. 2b) are ideal for printed applications due to their regularity and simplicity. They operate by comparing the input voltage to multiple reference voltages using an array of analog comparators. The comparator, i.e., the basic building block of the Flash ADC, is presented in Fig. 2a. This parallel, regular structure allows easy implementation. The comparators’ output represents the digital value in thermometer code, which is then converted to a binary encoding by counting the number of ‘1’s with a digital parallel counter. An $N$-bit ADC requires $2^{N}-1$ analog comparators.
SAR ADCs (Fig. 2c) sequentially compare the input signal to reference levels, adjusting bit-by-bit, from most to least significant bits. They require only one analog comparator, but also need a Digital-to-Analog Converter (DAC) and more complex digital logic (SAR logic) to store intermediate results and set reference values. Moreover, SAR ADCs require registers that are highly area-inefficient in EGFET. Hence, in SAR logic we use only an $N$-bit register for the output and a $\log_{2}N$ counter, with all other signals generated through counter multiplexing.
@@ III-B Arbitrary Precision Bespoke TNN Circuit
Focusing on area efficiency, a primary constraint in printed circuits [14], we use a single hidden layer with input precision ranging from $1$ to $4$ bits. Low-bit inputs allow for significantly smaller and less power-intensive ADCs, while also reducing the input precision requirements of the classifier (i.e., TNN). Throughout recent research efforts in ML circuits for printed electronics [16, 17, 27, 33], low input precision (i.e., below 4 bits) has been commonly used since target domains have low performance and precision needs, enabling printed circuits with acceptable area and power overheads [23].
@@
where $w_{i}$ are the neuron weights, $x_{i}$ are the input activations, and $n$ is the number of inputs. This reduces the precision required to represent $y$, with 1-1 being encoded as 0 while $1$ remains $1$. Note that (1) applies to any input precision, including $1$-bit inputs.
@@
where $w_{j}$ are the neurons’ weights and $y_{j}$ are the outputs of the hidden layer. However, considering the aforementioned encoding for $y_{j}$, negation is achieved with a single NOT gate ($0\rightarrow 1$ and $1\rightarrow 0$). In addition, due to the bespoke implementation, if a weight is 1-1, the corresponding activation is inverted before added to the respective popcount, whereas if the weight is 0, the corresponding connection is removed. The product of a zero weight by the respective activation is always 0 (numerical value). The (numerical) 0 is in the middle of 1-1 and $1$ and, thus, in our encoding 0 should be mapped to $\frac{1}{2}$ (i.e., $\frac{-1+1}{2}\rightarrow\frac{0+1}{2}$). Therefore, a $\frac{1}{2}$ is missing from the result for each removed connection due to zero weight. Hence, the output neuron can be implemented by a simple popcount unit (the sum in (2) accumulates only zeros or ones) and to obtain the correct output, a constant correction term $\frac{Z}{2}$ must be added to the popcount result, where $Z$ equals the number of zero weights in that neuron. Instead, we right-shift the popcount result $P$ and add $Z$: $2P+Z$. Note that $Z$ is a constant known post-training. Since the activation function of the output neuron is argmax, $2P+Z$ gives the same result as $P+\frac{Z}{2}$ without any rounding errors from integer division. As a result, the output neuron computes the following:
@@ IV-A2 LTG Error Analysis
If $\mathrm{LTG}_{\boldsymbol{w}}$ and $\mathrm{\hat{LTG}}_{\boldsymbol{w}}$ differ, it means that the two functions produced outputs of opposite signs–that is, one computed a positive weighted sum and the other a negative one. However, since $\mathrm{\hat{LTG}}_{\boldsymbol{w}}$ is implemented as a black-box function via CGP, we only observe its binary output (0 or $1$), without any direct insight into how closely it approximated the actual weighted sum. To capture the severity of such mismatches, we define an error metric $D(\boldsymbol{x})$ that quantifies the difference/distance between the weighted sum of $\mathrm{LTG}_{\boldsymbol{w}}$ and the decision threshold at 0, using the exact LTG as the reference value. This provides a meaningful measure of how far the approximate function was from producing the correct output, assuming that $\mathrm{\hat{LTG}}_{\boldsymbol{w}}$ flipped the sign of the sum. An offset of $1$ is added to handle cases where the sum is zero, but $\mathrm{\hat{LTG}}_{\boldsymbol{w}}$ wrongly produced a negative sign.
@@ V Results and Analysis
TNN Training: Exact TNN models are used as the baseline for accuracy comparisons. Input features are normalized within $[0,1]$, datasets follow a $70\%/30\%$ split for training/testing and TNNs are trained using QKeras [41] and its quantized_bits functionality. We employ the Adam optimizer for up to $30$ epochs of training (with early stopping). The learning rate for each network is obtained through Bayesian optimization, aiming to maximize the model’s inference accuracy with a maximum of $100$ attempts. Learning rate parameters are selected from the range of $0.001$ to $0.01$. Hidden neurons are selected via grid search within the $[1,50]$ range. For each dataset, we train a TNN for input precisions from $1$ to $4$ bits and, for each precision, select the TNN with the smallest hidden-layer size that achieves accuracy closest to the exact MLP [14].
@@ V-A1 Popcount Circuits
We generate approximate popcount circuits with the following constraints and configurations. The error limits $\tau_{mae}$ and $\tau_{wcae}$ are logarithmically distributed from $0.1$ to $0.5\cdot 2^{g}$ and from $1$ to $0.5\cdot 2^{m}$, with $g=\lceil\log_{2}m\rceil$, resulting in 2,0902,\!090 approximate circuits. We also set CGP search termination criteria to $30$, $60$, and $300$ minutes for popcount sizes $m<16$, $m<32$, and $m<60$, respectively. Using BDD evaluation, we achieve average speeds of 7,7597,\!759 and 1,3621,\!362 evaluations per second for $16$-bit and $32$-bit popcounts. For $60$-bit popcount circuits, this number falls to $25$ evaluations per second. To account for this, we extend the time limit, thereby increasing the chances of identifying solutions with better area-accuracy trade-offs.
@@
A visualization of the obtained results for three popcount circuit sizes is provided in Fig. 6. The featured circuits are synthesized and evaluated based on the error metrics and methodology described above. For easier comparison, area results are normalized relative to the exact popcount circuit of the same size. Our results indicate that as the error limit increases, CGP optimization identifies increasingly area-efficient solutions. Additionally, we assess the efficacy of our approximation approach through a comparison of our popcount circuits with variants derived using the truncation approximation approach, which has previously been applied in a wide range of approximate circuits, including MLPs [42]. As shown in Fig. 6, our approximation achieves significantly superior trade-offs compared to truncation, and delivers about 22x area reduction for $\varepsilon_{mae}$ of only $0.5$, $1.1$, and $1.9$ for $8$-bit, $16$-bit, and $47$-bit popcounts, respectively. Similar results to those shown in Fig. 6 are observed for all popcount circuit sizes.
@@ V-B Approximate TNN Evaluation
Fig. 9 shows the results of our multi-objective optimization for the $2$-bit Pendigits dataset using various approximate LTG libraries, each evolved with a distinct error metric. Libraries containing only Pareto-optimal LTG units (“pareto”) yield superior outcomes. For high accuracy, LTG designs optimized for the $\varepsilon_{wcde}$ metric are better, while those evolved for the $\varepsilon_{mde}$ metric are preferable for maximizing area gains. Similar findings to Fig. 9 apply to other datasets. Hence, for each TNN, the required LTG circuits are approximated against both $\varepsilon_{mde}$ and $\varepsilon_{wced}$ to populate our library. In Fig. 10, we further investigate the role of our $D$ error metric in approximating LTG units for achieving area-efficient and high-accuracy approximate TNNs. Fig. 10 presents a correlation analysis between $\varepsilon_{mde}$ and classification accuracy, when all TNN hidden neurons are replaced with approximate LTGs of identical $\varepsilon_{mde}$, for a wide range of error values. Two TNNs are used in this analysis: our most complex one (Pendigits) and a relatively small one (RedWine). A strong non-linear correlation between our $\varepsilon_{mde}$ metric and the classification accuracy can be observed, especially as the network’s knowledge capacity increases–i.e., when moving from smaller (RedWine) to larger (Pendigits) networks, or from lower ($1$-bit) to higher ($4$-bit) input precision. Fig. 10 validates that we obtain the expected correlation between neural network inference accuracy and approximate arithmetic units, and confirms that our $\varepsilon_{mde}$ metric is suitable for effectively guiding approximation exploration toward high-accuracy regions. Specifically, as $\varepsilon_{mde}$ (and thus area savings) increases, there exists a broad region where minimal accuracy degradation is observed at the TNN level, maximizing, thus, the area efficiency of our approximate TNNs.
@@
Fig. 11 shows the accuracy-area trade-off for all explored solutions during our multi-objective optimization, with the Pareto fronts (per input precision) annotated by lines. As Fig. 11 includes 15,00015,\!000 designs, synthesizing all is impractical, so we report the estimated area using our precise surrogate model. Post-synthesis results for key points are presented in Table II. Note that the area estimation in Fig. 11 reflects only the TNN classifier area, excluding any interfacing costs. In Fig. 11, we include only the input precisions that result in at least one Pareto-optimal point for each dataset, since precisions that lead to only dominated solutions (i.e., with higher area and lower accuracy) are suboptimal in our evolutionary-based optimization. As shown, our framework consistently delivers a smooth accuracy-area trade-off. Supporting arbitrary input precision allows us to effectively populate the accuracy-area Pareto front, as it is primarily composed of designs with varying precisions, with only one precision, especially the low cost $1$ bit, rarely dominating.
@@
Two cases require further discussion. For Pendigits with $1$-bit input precision, our framework maintains a smooth accuracy-area trade-off despite a significant initial accuracy loss w.r.t the exact $1$-bit TNN. Pendigits, being the most complex dataset examined, becomes highly sensitive to approximation with only $1$-bit inputs. This is evident from the $1$-bit exact Pendigits TNN achieving just 7676% accuracy, while the target MLP reaches 9494[14]. This further highlights that it is mandatory to support arbitrary input precision, as $1$-bit precision only (e.g., [20]) may be insufficient in many cases.
@@
For Arrhythmia, we present results only for $1$-bit precision, even though the $1$-bit exact TNN exhibits a 22% accuracy loss compared to the respective exact MLP, while the $2$-bit exact TNN features no accuracy loss. Our framework uses CGP to evolve exact LTG circuits into approximate variants and supports exact design with up to $90$ input bits. The $2$-bit Arrhythmia LTG circuits feature over $100$ inputs of $2$ bits, totaling more than $200$ input bits, and our CGP algorithm could not find suitable approximations within a reasonable timeframe. However, considering datasets like Arrhythmia, with more than $250$ input features, might be unrealistic for printed applications, particularly when factoring in ADC costs. Nevertheless, our framework still provides a strong trade-off, delivering 8686x lower area than the printed exact MLP [14], for only 55% accuracy loss (see Table II). In addition, our framework can be combined with pruning—which is not considered in this work to ensure fairness in comparisons—to reduce inputs to LTGs and potentially mitigate such issues. Alternatively, we can follow conventional hierarchical approaches for the generation of approximate LTGs with more than $200$ input bits. This involves breaking the LTG down to components (two or more adders and one comparator) and following a typical procedure for approximating dataflows [44, 45], e.g., creating a library of approximate components and then combining them through multi-objective optimization (e.g., NSGA-II). Such approaches might lead to some optimality loss but can address scalability issues in LTG approximation.